OpenStudy (anonymous):

What is the equation of the line, in point-slope form, that passes through the points (-1, 7) and (-2, 3)?

OpenStudy (anonymous):

find the slope using \[\frac{ y_2-y_1 }{ x_2-x_1 }\]formula then put that into point slope formula of \[y-y_1 = m(x-x_1) \]m is the slope

OpenStudy (anonymous):

can you give me the answer?

OpenStudy (anonymous):

can not give you the answer, but i can lead you to it it is part of the Code Of Conduct

OpenStudy (mathmale):

Rome, please find the slope of that line connecting the 2 given points.